# A question of Erdős and Graham on Egyptian fractions

### v1

@@ 1 Introduction
The study of Egyptian fractions, that is, sums of reciprocals of distinct positive integers, has a long history in combinatorial number theory (see, for example, [4]). The fact that every positive fraction can be written as an Egyptian fraction goes back at least to work of Fibonacci at the start of the 13th century. Much more recently, a result of Bloom [2] says that any subset of the natural numbers of positive upper density has a finite subset the sum of whose reciprocals adds to one.
In this paper, we will be concerned with a problem raised by Erdős and Graham [8, Page 36] in 1980 (see also [3, Problem 297]): how many ways are there to write one as a sum of distinct unit fractions with denominator at most $n$? Very recently, Steinerberger [11] showed that the number of such Egyptian fractions is at most $2^{0.93n}$. This already answered one particular question of Erdős and Graham, who asked whether the answer was $2^{n-o(n)}$. Here we answer their question much more precisely by showing that the count is $2^{(1+o_{n}(1))cn}$ for an explicit constant $c\approx 0.91117$.
@@
Our proof has two main steps. In the first step, we use entropy methods to show that $2^{(c_{x}+o(1))n}$ is the correct asymptotic count for the number of Egyptian fractions formed by adding distinct unit fractions with denominator at most $n$ whose sum is at most $x$. Then, in the second step, we use a method reminiscent of the absorption technique in extremal graph theory to show that the same asymptotic count holds for the number of Egyptian fractions summing to exactly $x$. Very roughly, we first set aside a small reservoir subset of $[n]$. Then, after finding many subsets of $[n]$ disjoint from this reservoir whose sums of reciprocals are somewhat smaller than $x$ and whose denominators have no very large prime power factors, we iteratively ‘clean’ these fractions by adding unit fractions from the reservoir to obtain a sum $x^{\prime}<x$ with small denominator. This is accomplished through the use of a recent result [5] on the existence of homogeneous generalized arithmetic progressions in subset sums. Finally, we find a small subset of the reservoir whose sum of reciprocals is equal to the remaining difference $x-x^{\prime}$.
@@ 2 Counting through entropy
In this section, we will use entropy methods to estimate the number of subsets $A\subseteq[n]$ with $s(A):=\sum_{a\in A}1/a\leq x$. To state our result, we need some notation. As in the introduction, let $h(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$ for $10<p<$ 1 and $h(0)=h(1)=0$. Given $x>0$, we choose $p_{1},\dots,p_{n}\in[0,1]$ so as to maximize $\sum_{m=1}^{n}h(p_{m})$ given $\sum_{m=1}^{n}p_{m}/m\leq x$. We then let $P(x)$ be the distribution of $(Y_{1},\dots,Y_{n})$, where each $Y_{m}=1$ with probability $p_{m}$ and 00 otherwise independently of each other. Then the (Shannon) entropy of $P(x)$ is given by $H(P(x))=\sum_{m=1}^{n}h(p_{m})$.
@@
Note that $H(P(x))$ is strongly convex and bounded (as a function of $p_{1},\dots,p_{n}$) in $[0,1]^{n}$. The unique stationary point of $H(P(x))$ in $[0,1]^{n}$ is $p_{1}=\dots=p_{n}=1/2$. Thus, for $x<(1/2)(\sum_{m=1}^{n}1/m)$, the maxima of $H(P(x))$ must be achieved on the boundary of $[0,1]^{n}\cap\{\sum_{m}p_{m}/m\leq x\}$ and, since $h(x)=0$ for $x\in\{0,1\}$, we must have that the maxima are achieved on $\sum_{m}p_{m}/m=x$. At a maximum $(p_{1}^{*},\dots,p_{n}^{*})$, we must have that $\nabla H(P(x))$ is parallel to $(1,1/2,\dots,1/n)$. Noting that $h^{\prime}(p)=\log\frac{1-p}{p}$, we obtain that such a point satisfies $p_{m}^{*}=\frac{1}{1+e^{c^{\prime}/m}}$ for some $c^{\prime}$. By the condition $\sum_{m\leq n}p_{m}^{*}/m=x$, we must also have that $c^{\prime}$ satisfies
@@
It is easy to check that $c>0$ when $x<(1-\varepsilon)(\ln n)/2$ and, for $x$ sufficiently large, that $c=\Theta(e^{-2x})$. Indeed, for the last estimate, we observe that
@@
[…]
@@
Let $X_{1},\dots,X_{n}$ be independent centered random variables with $\mathbb{E}[X_{i}^{2}]=\zeta_{i}$ and $\mathbb{E}[X_{i}^{3}]=\rho_{i}$. Let $Z=\frac{\sum_{i=1}^{n}X_{i}}{(\sum_{i=1}^{n}\zeta_{i})^{1/2}}$ and $Z^{\prime}$ be a standard Gaussian. Then
@@
To see the last bound, observe that $\sum_{m=cn/(k+1)}^{cn/k}\frac{e^{-cn/m}}{m^{2}}=\Theta(\frac{e^{-k}}{cn})$. Similarly, for $1\leq k\leq 1/c$, $\sum_{m=cnk}^{cn(k+1)}\frac{e^{-cn/m}}{m^{2}}=\Theta(\frac{e^{-1/k}}{cnk^{2}})$. The desired bound follows from summing these estimates over $k$.
By a similar argument, the sum of the centered third moments of the $Y_{i}$ is
$\sum_{m=1}^{n}\left[\frac{1}{1+e^{cn/m}}\left(\frac{1}{m}-\frac{1}{m(1+e^{cn/m})}\right)^{3}+\left(1-\frac{1}{1+e^{cn/m}}\right)\left(-\frac{1}{m(1+e^{cn/m})}\right)^{3}\right]$ $=O\left(\sum_{m=1}^{n}\frac{e^{-cn/m}}{m^{3}}\right)$
@@
and we will use a similar argument with the Berry–Esseen bound to show that $\mathrm{Pr}(Y_{m}=1|E=0)$ is close to $\mathrm{Pr}(Y_{m}=1)$. Indeed, the calculations above similarly yield that the random variable $Z^{\prime}_{m}=Z-Y_{m}+1/m$ is a sum of independent random variables with $\mathbb{E}Z^{\prime}_{m}=x+\frac{1}{m}-\frac{1}{m(1+e^{cn/m})}$, $\mathrm{Var}(Z^{\prime}_{m})=\Theta(1/(cn))$ and the sum of centered third moments $O(1/(cn)^{2})$. By Lemma 4, for $g\sim\mathcal{N}(0,1)$,
@@
where in the last inequality we used that $\mathrm{Pr}(Y_{m}=1)\leq 1/2$, so $\log\frac{1}{\mathrm{Pr}(Y_{m}=1)\left(1-O\left(\frac{\sqrt{cn}}{m}+\frac{1}{\sqrt{cn}}\right)\right)}=O\left(\log\frac{1}{\mathrm{Pr}(Y_{m}=1)}\right)$. Therefore,
@@
Again, for the last estimate, we note that $\sum_{m=cn/(k+1)}^{cn/k}\left(\frac{\sqrt{cn}}{m}+\frac{1}{\sqrt{cn}}\right)\frac{cn/m}{1+e^{cn/m}}\leq O\left(\frac{cn}{k^{2}}e^{-k}\frac{k^{2}}{\sqrt{cn}}\right)=O(e^{-k}\sqrt{cn})$ and, for $1\leq k\leq 1/c$, $\sum_{m=cnk}^{cn(k+1)}\left(\frac{\sqrt{cn}}{m}+\frac{1}{\sqrt{cn}}\right)\frac{cn/m}{1+e^{cn/m}}\leq O\left(cn\frac{1}{k\sqrt{cn}}\right)\leq O\left(\frac{1}{k}\sqrt{cn}\right)$. Summing over $k$, we thus have
@@ 3 Subset sums of modular inverses
The main technical tool we still need is the following result, which says that if $q$ is a large prime power and we take a dense subset $I$ of the interval $[q^{\varepsilon},2q^{\varepsilon}]$, then every residue class mod $q$ can be written as the sum of a small number of reciprocals of elements of $I$. Roughly speaking, this allows us to cancel out any particular prime power from the denominator of a fraction in the absorption step of the proof of Theorem 1. Given a set $A$ of integers, we will use the notation $\Sigma^{[s]}(A)$ for the collection of sums of subsets of $A$ of size at most $s$.
@@
Let $\delta,\varepsilon>0$ and let $q$ be a prime power which is sufficiently large in terms of $\delta,\varepsilon$. If $I$ is a subset of $[q^{\varepsilon},2q^{\varepsilon}]$ consisting of elements coprime to $q$ with $|I|\geq\delta q^{\varepsilon}$, then $\Sigma^{[s]}(I^{-1})\pmod{q}=\mathbb{Z}_{q}$ for $s=q^{\varepsilon/2}$.
In the proof of Theorem 5, we will make use of the following key result from [5]. Recall that a generalized arithmetic progression (henceforth GAP) $P$ of dimension $k$ is a set of integers $\{x_{0}+\ell_{1}x_{1}+\ell_{2}x_{2}+\cdots+\ell_{k}x_{k}|0\leq\ell_{1}<L_{1},\ldots,0\leq\ell_{d}<L_{k}\}$. A GAP is called proper if it has size exactly $L_{1}L_{2}\cdots L_{k}$. We say that $P$ is homogeneous if $x_{0}$ divides $x_{1},\dots,x_{k}$. For a natural number $t$, we define $tP$ to be the $t$-fold sumset of $P$, while if $t$ is a positive real number which is not an integer and $P=\{\sum_{i=1}^{k}n_{i}x_{i}:a_{i}\leq n_{i}\leq b_{i}\}$ is a homogeneous GAP, we can generalize the definition by setting $tP=\{\sum_{i=1}^{k}n_{i}x_{i}:ta_{i}\leq n_{i}\leq tb_{i}\}$.
@@
For any $\beta>1$ and $10<\eta<$ 1, there are positive constants $c$ and $k$ such that the following holds. Let $A$ be a subset of $[n]$ of size $m$ with $n\leq m^{\beta}$ and let $s\in[m^{\eta},cm/\log m]$. Then there exists a subset $\hat{A}$ of $A$ of size at least $m-c^{-1}s\log m$ and a proper GAP $P$ of dimension at most $k$ such that $\hat{A}\cup\{0\}$ is a subset of $P$. Furthermore, there exists $A^{\prime}\subseteq\hat{A}$ of size at most $s$ such that $\Sigma(A^{\prime})$ contains a homogeneous translate of $csP$, where $csP$ is proper.
We will also need the following simple variant of Dirichlet’s simultaneous approximation theorem. For a residue class $i\pmod{q}$, we use the notation $\overline{i}$ for the unique integer in $(-q/2,q/2]$ congruent to $i$ modulo $q$.
@@
Let $s=q^{\varepsilon/2}$. Let $\overline{I}$ and $\overline{I^{-1}}$ denote the set of integer representations (in $(-q/2,q/2]$) of $I$ and $I^{-1}$. By Theorem 6, there is $c$ depending only on $\varepsilon$ such that we can find $J\subseteq\overline{I^{-1}}$ of size at least $|I|-c^{-1}s\log|I|=(1-o(1))|I|$ and a proper GAP $P$ of dimension $k=O_{\varepsilon}(1)$ such that $J\cup\{0\}\subseteq P$ and $\Sigma^{[s]}(J)$ contains a translate of $csP$ which is proper.
@@
On the other hand, since $\Sigma^{[s]}(J)$ contains a translate of $csP$ with $csP$ proper and $\Sigma^{[s]}(J)\subseteq(-sq/2,sq]$, we have that
@@ 4 Absorption
Let $\varepsilon>0$ be sufficiently small. Then there exists $\xi>0$ such that if $x\leq\xi\ln n$ is a rational whose denominator is $(n^{1-\varepsilon}/2)$-powersmooth, then the number of subsets $A\subseteq[n]$ with $x=\sum_{a\in A}1/a$ is at least $2^{c_{x}n-c_{\varepsilon}n}$, where $c_{\varepsilon}\to 0$ as $\varepsilon\to 0$ and
[…]
@@
Hence, at least $(1-\frac{3}{2}\delta)n-o(n)\geq(1-2\delta)n$ positive integers at most $n$ are $t$-powersmooth, as required. ∎
@@
By choosing $c_{\varepsilon}$ suitably, we can assume that $n$ is sufficiently large in terms of $\varepsilon$. Let $L$ be sufficiently large, assuming in particular that Theorem 5 applies for $\varepsilon$ as in the statement of the theorem, $\delta=\frac{1}{2}$ and all $q>L$. Let $K$ denote the least common multiple of all prime powers at most $L$. We first reserve the set $\mathcal{R}$ of multiples of $K$ in $[n]$. Let $\mathcal{P}(q)=q\cdot[q^{\varepsilon},2q^{\varepsilon}]\setminus\mathcal{R}$ and $\mathcal{P}=\bigcup_{q}\mathcal{P}(q)$, where $q$ ranges over all prime powers at most $n^{1-\varepsilon}/2$. Here $q\cdot S=\{qs:s\in S\}$ and the notation $[q^{\varepsilon},2q^{\varepsilon}]$ refers to the set of integers in this interval. Let $\mathcal{S}$ denote the set of $(n^{1-\varepsilon}/2)$-powersmooth numbers at most $n$ and $\mathcal{U}=\mathcal{S}\setminus(\mathcal{R}\cup\mathcal{P})$. Lemma 9 implies that $|\mathcal{S}|\geq(1-O(\varepsilon))n$ and we also have that $|\mathcal{P}\cap[n]|\leq\sum_{q\leq n^{1-\varepsilon}}2q^{\varepsilon}\leq 4\frac{n}{\log n}$. Thus,
@@
Fix one such sum corresponding to a set $A_{0}\subseteq[n]$, and let $x_{0}=x-s(A_{0})\geq\eta x$. Consider the following procedure, where at each step $i$ we have a real number $x_{i}$ and a set $A_{i}$ for which $s(A_{i})=x_{i}$:
@@
    In decreasing order over the prime powers larger than $L$, consider the largest prime power $q=q_{i}\leq n^{1-\varepsilon}/2$ of a prime $p=p_{i}$ which appears as a factor of the denominator of $x_{i}$. We then find $B_{i}\subseteq(1/q)\cdot\mathcal{P}(q)$ of size at most $q^{\varepsilon/2}$ such that, for $x_{i}=\frac{u_{i}}{v_{i}}$ with $u_{i},v_{i}$ coprime, $s(B_{i})=-\frac{u_{i}}{v_{i}/q}\pmod{q}$. We say that step $i$ succeeds if we can find such a $B_{i}$. If it does succeed, we update $x_{i+1}=x_{i}-s(q\cdot B_{i})$ and $A_{i+1}=A_{i}\cup q\cdot B_{i}$, noting by our choice that no nonzero power of $p$ divides the denominator of $x_{i+1}$. Furthermore, any new prime power divisor of $x_{i+1}$ is at most $2q^{\varepsilon}$.
@@
We ensure that $L,\eta>0$ are chosen (depending on $\varepsilon$) so that $C\varepsilon^{-1}L^{-\varepsilon/2}<\eta x/2$. From (9), we have that $x\geq x_{f}$ and $x-x_{f}$ is a positive rational number at most $x$ whose denominator is $L$-powersmooth. As such, we have that $K(x-x_{f})$ is a positive integer with $K(x-x_{f})\leq Kx$. We now note that there exists a subset $D\subseteq[n/K]$ such that $s(D)=K(x-x_{f})$, where we use the assumption that $x\leq\xi\ln n$ for $\xi$ chosen sufficiently small in $\varepsilon$, so that $Kx<\varepsilon\log(n/K)$. To see that this is the case, one may, for example, make use of Croot’s result [6] that one can always be written as the sum of reciprocals of numbers from any interval of the form $[t,(e+o(1))t]$. This allows us to iteratively remove $K(x-x_{f})$ disjoint subsets from $[n/K]$, the sum of the reciprocals of each of which is one. We then set their union to be $D$, noting that $K\cdot D\subseteq\mathcal{R}$ is disjoint from $\mathcal{U}$ and $\mathcal{P}$.
@@
Note added. As we completed this paper, we learned that a result similar to our Theorem 1 was obtained simultaneously and independently, though using rather different methods, by Yang P. Liu and Mehtaab Sawhney [9].
Acknowledgements. We are grateful to the American Institute of Mathematics for hosting the SQuaREs project at which this work was initiated. Research supported by NSF Awards DMS-2054452 and DMS-2348859 (David Conlon), NSF Award DMS-2154129 (Jacob Fox), NSF Award DMS-2103154 (Xiaoyu He), NSF Awards DMS-1952767 and DMS-2153576 (Dhruv Mubayi), a Clay Research Fellowship and a Stanford Science Fellowship (Huy Tuan Pham), an NSF CAREER Award and NSF Awards DMS-1952786 and DMS-2246847 (Andrew Suk) and NSF Award DMS-1800332 (Jacques Verstraëte).

### v2

@@ 1 Introduction
The study of Egyptian fractions, that is, sums of reciprocals of distinct positive integers, has a long history in combinatorial number theory (see, for example, [4]). The fact that every positive fraction can be written as an Egyptian fraction goes back at least to work of Fibonacci at the start of the 13th century. Much more recently, a result of Bloom [2] says that any subset of the natural numbers of positive upper density has a finite subset the sum of whose reciprocals adds to 1.
In this paper, we will be concerned with a problem raised by Erdős and Graham [8, Page 36] in 1980 (see also [3, Problem 297]): how many ways are there to write 1 as a sum of distinct unit fractions with denominator at most $n$? In 2017, this problem was partially addressed by several users, including Lucia, RaphaelB4 and js21, in response to a MathOverflow post [11] by Mikhail Tikhomirov. They showed that there is an explicit constant $c\approx 0.91117$ such that the number of such Egyptian fractions is at most $2^{(1+o_{n}(1))cn}$. This already answered one particular question of Erdős and Graham, who asked whether the answer was $2^{n-o(n)}$. Here we solve Erdős and Graham’s problem more precisely by showing that the count is equal to $2^{(1+o_{n}(1))cn}$ for the same $c\approx 0.91117$.
@@
Our proof has two main steps. In the first step, we use entropy methods to show that $2^{(c_{x}+o(1))n}$ is the correct asymptotic count for the number of Egyptian fractions formed by adding distinct unit fractions with denominator at most $n$ whose sum is at most $x$. This may also be obtained using the techniques discussed in the MathOverflow post [11], but we include our proof, which we found before being alerted to that post, for completeness. Then, in the second step, we use a method reminiscent of the absorption technique in extremal graph theory to show that the same asymptotic count holds for the number of Egyptian fractions summing to exactly $x$. Very roughly, we first set aside a small reservoir subset of $[n]$. Then, after finding many subsets of $[n]$ disjoint from this reservoir whose sums of reciprocals are somewhat smaller than $x$ and whose elements have no very large prime power factors, we iteratively ‘clean’ these fractions by adding unit fractions from the reservoir to obtain a sum $x^{\prime}<x$ with small denominator. This is accomplished through the use of a recent result [5] on the existence of homogeneous generalized arithmetic progressions in subset sums. Finally, we find a small subset of the remaining reservoir whose sum of reciprocals is equal to the remaining difference $x-x^{\prime}$.
@@ 2 Counting through entropy
In this section, we will use entropy methods to estimate the number of subsets $A\subseteq[n]$ with $s(A):=\sum_{a\in A}1/a\leq x$. To state our result, we need some notation. As in the introduction, let $h(p)=-p\log_{2}p-(1-p)\log_{2}(1-p)$ for $0<p<1$ and $h(0)=h(1)=0$. Given $x>0$, we choose $p_{1},\dots,p_{n}\in[0,1]$ so as to maximize $\sum_{m=1}^{n}h(p_{m})$ given $\sum_{m=1}^{n}p_{m}/m\leq x$. We then let $P(x)$ be the distribution of $(Y_{1},\dots,Y_{n})$, where each $Y_{m}=1$ with probability $p_{m}$ and 0 otherwise independently of each other. Then the (Shannon) entropy of $P(x)$ is given by $H(P(x))=\sum_{m=1}^{n}h(p_{m})$.
@@
Note that $H(P(x))$ is strongly convex and bounded (as a function of $p_{1},\dots,p_{n}$) in $0,1]^{n}$. The unique stationary point of $H(P(x))$ in $0,1]^{n}$ is $p_{1}=\dots=p_{n}=1/2$. Thus, for $x<(1/2)(\sum_{m=1}^{n}1/m)$, the maxima of $H(P(x))$ must be achieved on the boundary of $0,1]^{n}\cap\{\sum_{m}p_{m}/m\leq x\}$ and, since $h(x)=0$ for $x\in\{0,1\}$, we must have that the maxima are achieved on $\sum_{m}p_{m}/m=x$. At a maximum $(p_{1}^{*},\dots,p_{n}^{*})$, we must have that $\nabla H(P(x))$ is parallel to $(1,1/2,\dots,1/n)$. Noting that $h^{\prime}(p)=\log\frac{1-p}{p}$, we obtain that such a point satisfies $p_{m}^{*}=\frac{1}{1+e^{c^{\prime}/m}}$ for some $c^{\prime}$. By the condition $\sum_{m\leq n}p_{m}^{*}/m=x$, we must also have that $c^{\prime}$ satisfies
@@
It is easy to check that $c>0$ when $x<(1-\varepsilon)(\ln n)/2$ and, for $x$ sufficiently large, that $c=\Theta(e^{-2x})$. Indeed, for the last estimate, we observe that, for $k\geq 1$,
@@
Finally, for $x\in[\varepsilon,1/\varepsilon]$ and $c>0$, we can approximate $\frac{1}{n}\sum_{m=1}^{n}\frac{1}{(m/n)(1+e^{c/(m/n)})}$ by the integral $\int_{0}^{1}\frac{1}{y(1+e^{c/y})}dy$. We thus obtain that for $\lambda>0$ satisfying $\int_{0}^{1}\frac{1}{y(1+e^{\lambda/y})}dy=x$, we have $c_{x,n}=\lambda+o_{n}(1)$, from which (1) follows readily. ∎
@@
Let $X_{1},\dots,X_{n}$ be independent centered random variables with $\mathbb{E}[X_{i}^{2}]=\zeta_{i}$ and $\mathbb{E}[|X_{i}|^{3}]=\rho_{i}$. Let $Z=\frac{\sum_{i=1}^{n}X_{i}}{(\sum_{i=1}^{n}\zeta_{i})^{1/2}}$ and $Z^{\prime}$ be a standard Gaussian. Then
@@
To see the last bound, observe that, for $k\geq 1$, $\sum_{m=cn/(k+1)}^{cn/k}\frac{e^{-cn/m}}{m^{2}}=\Theta(\frac{e^{-k}}{cn})$. Similarly, for $1\leq k\leq 1/c$, $\sum_{m=cnk}^{cn(k+1)}\frac{e^{-cn/m}}{m^{2}}=\Theta(\frac{e^{-1/k}}{cnk^{2}})$. The desired bound follows from summing these estimates over $k$.
By a similar argument, the sum of the absolute centered third moments of the $Y_{m}/m$ is
$\sum_{m=1}^{n}\left[\frac{1}{1+e^{cn/m}}\left|\frac{1}{m}-\frac{1}{m(1+e^{cn/m})}\right|^{3}+\left(1-\frac{1}{1+e^{cn/m}}\right)\left|-\frac{1}{m(1+e^{cn/m})}\right|^{3}\right]$ $=O\left(\sum_{m=1}^{n}\frac{e^{-cn/m}}{m^{3}}\right)$
@@
and we will use a similar argument with the Berry–Esseen bound to show that $\mathrm{Pr}(Y_{m}=1|E=0)$ is close to $\mathrm{Pr}(Y_{m}=1)$. Indeed, the calculations above similarly yield that the random variable $Z^{\prime}_{m}=Z-Y_{m}/m+1/m$ is a sum of independent random variables with $\mathbb{E}Z^{\prime}_{m}=x+\frac{1}{m}-\frac{1}{m(1+e^{cn/m})}$, $\mathrm{Var}(Z^{\prime}_{m})=\Theta(1/(cn))$ and the sum of absolute centered third moments $O(1/(cn)^{2})$. By Lemma 3, for $g\sim\mathcal{N}(0,1)$,
@@
where in the last inequality we used $\mathrm{Pr}(Y_{m}=1)\leq 1/2$, so $\log\frac{1}{\mathrm{Pr}(Y_{m}=1)\left(1-O\left(\frac{\sqrt{cn}}{m}+\frac{1}{\sqrt{cn}}\right)\right)}=O\left(\log\frac{1}{\mathrm{Pr}(Y_{m}=1)}\right)$. Therefore,
@@
Again, for the last estimate, we note that, for $k\geq 1$, $\sum_{m=cn/(k+1)}^{cn/k}\left(\frac{\sqrt{cn}}{m}+\frac{1}{\sqrt{cn}}\right)\frac{cn/m}{1+e^{cn/m}}\leq O\left(\frac{cn}{k^{2}}e^{-k}\frac{k^{2}}{\sqrt{cn}}\right)=O(e^{-k}\sqrt{cn})$ and, for $1\leq k\leq 1/c$, $\sum_{m=cnk}^{cn(k+1)}\left(\frac{\sqrt{cn}}{m}+\frac{1}{\sqrt{cn}}\right)\frac{cn/m}{1+e^{cn/m}}\leq O\left(cn\frac{1}{k\sqrt{cn}}\right)\leq O\left(\frac{1}{k}\sqrt{cn}\right)$. Summing over $k$, we thus have
@@ 3 Subset sums of modular inverses
The main technical tool we still need is the following result, which says that if $q$ is a large prime power and we take a dense subset $I$ of the interval $q^{\varepsilon},2q^{\varepsilon}]$, then every residue class mod $q$ can be written as the sum of a small number of reciprocals of elements of $I$. Roughly speaking, this allows us to cancel out any particular prime power from the denominator of a fraction in the absorption step of the proof of Theorem 1. Given a set $A$ of integers, we will use the notation $\Sigma^{[s]}(A)$ for the collection of sums of subsets of $A$ of size at most $s$.
@@
Let $\delta,\varepsilon>0$ and let $q$ be a prime power which is sufficiently large in terms of $\delta,\varepsilon$. If $I$ is a subset of $q^{\varepsilon},2q^{\varepsilon}]$ consisting of elements coprime to $q$ with $|I|\geq\delta q^{\varepsilon}$, then $\Sigma^{[s]}(I^{-1})\pmod{q}=\mathbb{Z}_{q}$ for $s=q^{\varepsilon/2}$.
In the proof of Theorem 2, we will make use of the following key result from [5]. Recall that a generalized arithmetic progression (henceforth GAP) $P$ of dimension $k$ is a set of integers $\{x_{0}+\ell_{1}x_{1}+\ell_{2}x_{2}+\cdots+\ell_{k}x_{k}|0\leq\ell_{1}<L_{1},\ldots,0\leq\ell_{k}<L_{k}\}$. A GAP is called proper if it has size exactly $L_{1}L_{2}\cdots L_{k}$. We say that $P$ is homogeneous if $\gcd(x_{1},\dots,x_{k})$ divides $x_{0}$. For a natural number $t$, we define $tP$ to be the $t$-fold sumset of $P$, while if $t$ is a positive real number which is not an integer and $P=\{\sum_{i=1}^{k}n_{i}x_{i}:a_{i}\leq n_{i}\leq b_{i}\}$ is a homogeneous GAP, we can generalize the definition by setting $tP=\{\sum_{i=1}^{k}n_{i}x_{i}:ta_{i}\leq n_{i}\leq tb_{i}\}$.
@@
[…]
We will also need the following simple variant of Dirichlet’s simultaneous approximation theorem. For a residue class $i\pmod{q}$, we use the notation $\overline{i}$ for the unique integer in (q/2,q/2](-q/2,q/2] congruent to $i$ modulo $q$.
@@
Let $s=q^{\varepsilon/2}$. Let $\overline{I}$ and $\overline{I^{-1}}$ denote the set of integer representations (in (q/2,q/2](-q/2,q/2]) of $I$ and $I^{-1}$. By Theorem 3, there is $c$ depending only on $\varepsilon$ such that we can find $J\subseteq\overline{I^{-1}}$ of size at least $|I|-c^{-1}s\log|I|=(1-o(1))|I|$ and a proper GAP $P$ of dimension $k=O_{\varepsilon}(1)$ such that $J\cup\{0\}\subseteq P$ and $\Sigma^{[s]}(J)$ contains a translate of $csP$ which is proper.
@@
On the other hand, since $\Sigma^{[s]}(J)$ contains a translate of $csP$ with $csP$ proper and $\Sigma^{[s]}(J)\subseteq(-sq/2,sq/2]$, we have that
@@ 4 Absorption
Let $\varepsilon>0$ be sufficiently small. Then there exists $\xi>0$ such that if $\varepsilon\leq x\leq\xi\ln n$ is a rational whose denominator is $(n^{1-\varepsilon}/2)$-powersmooth, then the number of subsets $A\subseteq[n]$ with $x=\sum_{a\in A}1/a$ is at least $2^{c_{x}n-c_{\varepsilon}n}$, where $c_{\varepsilon}\to 0$ as $\varepsilon\to 0$ and
$c_{x}:=\int_{0}^{1}h\left(\frac{1}{1+e^{\lambda/y}}\right)dy$
@@
Hence, at least $(1-\frac{3}{2}\delta)n-o(n)\geq(1-2\delta)n$ positive integers at most $n$ are $t$-powersmooth. ∎
@@
By choosing $\xi$ suitably, we can assume that $n$ is sufficiently large in terms of $\varepsilon$. Let $L$ be sufficiently large, assuming in particular that Theorem 2 applies for $\varepsilon$ as in the statement of the theorem, $\delta=\frac{1}{2}$ and all $q>L$. Let $K$ denote the least common multiple of all prime powers at most $L$. We first reserve the set $\mathcal{R}$ of multiples of $K$ in $[n]$. Let $\mathcal{P}(q)=q\cdot[q^{\varepsilon},2q^{\varepsilon}]\setminus\mathcal{R}$ and $\mathcal{P}=\bigcup_{q}\mathcal{P}(q)$, where $q$ ranges over all prime powers at most $n^{1-\varepsilon}/2$. Here $q\cdot S=\{qs:s\in S\}$ and the notation $q^{\varepsilon},2q^{\varepsilon}]$ refers to the set of integers in this interval. Let $\mathcal{S}$ denote the set of $(n^{1-\varepsilon}/2)$-powersmooth numbers at most $n$ and $\mathcal{U}=\mathcal{S}\setminus(\mathcal{R}\cup\mathcal{P})$. Lemma 5 implies that $|\mathcal{S}|\geq(1-O(\varepsilon))n$ and we also have that $|\mathcal{P}\cap[n]|\leq\sum_{q\leq n^{1-\varepsilon}}2q^{\varepsilon}\leq 4\frac{n}{\ln n}$. Thus,
@@
Fix one such sum corresponding to a set $A_{0}\subseteq[n]$ and let $x_{0}=x-s(A_{0})\geq\eta x$. Consider the following procedure, where at each step $i$ we have a real number $x_{i}$ and a set $A_{i}$ for which $x_{i}=x-s(A_{i})$:
@@
    In decreasing order over the prime powers larger than $L$, consider the largest prime power $q=q_{i}\leq n^{1-\varepsilon}/2$ of a prime $p=p_{i}$ which appears as a factor of the denominator of $x_{i}$. We then find $B_{i}\subseteq(1/q)\cdot\mathcal{P}(q)$ of size at most $q^{\varepsilon/2}$ such that, for $x_{i}=\frac{u_{i}}{v_{i}}$ with $u_{i},v_{i}$ coprime, $s(B_{i})=-\frac{u_{i}}{v_{i}/q}\pmod{q}$. We say that step $i$ succeeds if we can find such a $B_{i}$. If it does succeed, we update $x_{i+1}=x_{i}-s(q\cdot B_{i})$ and $A_{i+1}=A_{i}\cup q\cdot B_{i}$, noting by our choice that no nonzero power of $p$ divides the denominator of $x_{i+1}$. Furthermore, any new prime power divisor of the denominator of $x_{i+1}$ is at most $2q^{\varepsilon}$.
@@
We ensure that $L,\eta>0$ are chosen (depending on $\varepsilon$) so that $C\varepsilon^{-1}L^{-\varepsilon/2}<\eta x/2$. From (9), we have that $x_{f}$ is a positive rational number at most $x$ whose denominator is $L$-powersmooth. As such, we have that $Kx_{f}$ is a positive integer with $Kx_{f}\leq Kx$. We now note that there exists a subset $D\subseteq[n/K]$ such that $s(D)=Kx_{f}$, where we use the assumption that $x\leq\xi\ln n$ for $\xi$ chosen sufficiently small in $\varepsilon$, so that $Kx<\varepsilon\ln(n/K)$. To see that this is the case, one may, for example, make use of Croot’s result [6] that 1 can always be written as the sum of reciprocals of numbers from any interval of the form $[t,(e+o(1))t]$. This allows us to iteratively remove $Kx_{f}$ disjoint subsets from $[n/K]$, the sum of the reciprocals of each of which is 1. We then set their union to be $D$, noting that $K\cdot D\subseteq\mathcal{R}$ is disjoint from $\mathcal{U}$ and $\mathcal{P}$.
@@
Note added
As we completed this paper, we learned that a result similar to our Theorem 1 was obtained simultaneously and independently, though using rather different methods, by Yang P. Liu and Mehtaab Sawhney [9].
Acknowledgments
We are grateful to the American Institute of Mathematics for hosting the SQuaREs project at which this work was initiated. We are also indebted to Zachary Chase and the user Lucia for bringing the MathOverflow post [11] to our attention.